The sum of two numbers is 2500. If 1/3 of the number is 1/2 of the other, find the numbers
Answer by b_gulsum(6)

a+b=2500
a/3=b/2
Then, a=3k and b=2k
5k=2500, k=500
a=1500 and b=1000